OpenStudy (anonymous):

An arc AB of a circle with center O & radius 20cm subtends an angle of 26.25 degrees at center. Calculate the length of the arc AB & the area of the sector bounded by the arc AB & the two radii OA & OB . . ANSWER: 9.16cm, 91.6cm^2

OpenStudy (anonymous):

no idea how to do this one :/

OpenStudy (anonymous):

\[arc=\theta \times radius\] theta must be in radian so angle 26.25 degree = 26.25*pi/180 radian \[arc=0.4579\times 20\] \[arc= 9.16 cm\] \[area of sector=\Pi \times r^2 \times \theta( degree)/360\] \[area of sector=\Pi \times 20^2 \times 26.25/360\] \[area of sector AOB=91.629 cm^2\]